can you give me step by step
-5x-4y=-89/5
4x-4y=46/5
for x, for y
thank you

Answer by xcentaur(357)

-5x-4y=-89/5
4x-4y=46/5


Solved by pluggable solver: Linear System solver (using determinant)
Solve:


Any system of equations:


has solution

or



(x=3, y=0.7}


Thus,(x=3, y=0.7}


Prabhat.